### v1

# Kolmogorov–Arnold Networks in Fraud Detection: Bridging the Gap Between Theory and Practice

Abstract
Kolmogorov–Arnold Networks (KAN) are highly efficient in inference and can handle complex patterns once trained, making them desirable for production environments and ensuring a fast service experience in the finance and electronic shopping industries. However, we found that KAN, in general, is not suitable for fraud detection problems. We also discovered a quick method to determine whether a problem is solvable by KAN: if the data can be effectively separated using spline interpolation with varying intervals after applying Principal Component Analysis (PCA) to reduce the data dimensions to two, KAN can outperform most machine learning algorithms. Otherwise, it indicates KAN may not solve the problem effectively compared to other machine learning algorithms. We also propose a heuristic approach for selecting the appropriate hyperparameters for KAN to significantly accelerate training time compared to grid search hyperparameter tuning, which usually takes a month for a comprehensive grid search. Specifically, the width parameter should generally follow a pyramid structure, allowing efficient spline mixing, and $k$ should be fixed at 15, with the grid number fixed at 5. This streamlined approach minimizes the number of evaluations required, significantly speeding up the hyperparameter tuning process while still achieving robust performance metrics.
1 Introduction
Fraud detection is a critical challenge in various industries, particularly finance, insurance, and e-commerce. Traditional machine learning techniques, such as logistic regression, decision trees, and ensemble methods, have been extensively employed to address this problem. These methods, while effective and widely adopted, may struggle to capture complex, non-linear relationships in the data. Kolmogorov–Arnold Networks (KAN) (Liu et al., 2024) offer a promising alternative due to their capability to represent intricate functions through a relatively simple structure. This feature makes KAN suitable for fraud detection applications where accurate and low false-positive decision-making is paramount.
However, our research indicates that KANs are not universally applicable for fraud detection tasks. One significant limitation is the considerable time required for training. To address this, we introduce a heuristic approach for selecting KAN hyperparameters, significantly reducing training time compared to the exhaustive grid search method, which demands extensive computational resources. Specifically, we recommend a pyramid structure for the width parameter, fixing $k$ at 15, and the grid number at 5. This strategy minimizes complexity from the brute force search’s $O(n^{3})$ to $O(1)$ and greatly enhances training efficiency for KAN.
Furthermore, KAN may not be the best solution for all types of fraud detection tasks. We propose a rapid assessment method to determine KAN’s suitability for a given problem: if the data, after dimensionality reduction using Principal Component Analysis (PCA) to two dimensions, can be effectively separated using spline interpolation with varying intervals, KAN can outperform most machine learning algorithms. Otherwise, it might not be worth the time to train KAN models, as other machine learning models can achieve similar or even better results.
By implementing these strategies, we aim to demonstrate both the potential and limitations of KAN in fraud detection, offering practical guidance for its application in real-world scenarios.
2 Related Work
Previous studies on fraud detection have predominantly employed traditional machine learning techniques such as logistic regression, decision trees, and ensemble methods. While these techniques have shown efficacy, they often require extensive feature engineering and may not capture complex, non-linear relationships in the data.
2.1 Multi-layer Perceptron
Deep learning approaches, such as Multi-Layer Perceptron (MLP) (Rosenblatt, 1958), have been explored to address these limitations. MLPs serve as foundational building blocks in deep learning due to their ability to model complex patterns. However, MLPs use the same activation functions across their layers, which requires a large amount of data to effectively represent the desired patterns. In contrast, Kolmogorov–Arnold Networks (KAN) offer flexible activation functions for each edge, making them more efficient in inference and pattern representation, thus presenting a potential alternative.
2.2 Dimensionality Reduction
Dimensionality reduction techniques have also been used to preprocess high-dimensional data, simplifying the model training process and improving performance. For instance, Jolliffe (2002) discusses Principal Component Analysis (PCA), which reduces dimensionality by transforming the data into a set of linearly uncorrelated components. PCA is computationally efficient and retains most of the variance in the data, making it suitable for our purposes.
Other techniques like Linear Discriminant Analysis (LDA) (Fisher, 1936), t-Distributed Stochastic Neighbor Embedding (t-SNE) (van der Maaten and Hinton, 2008), and Autoencoders (Hinton and Salakhutdinov, 2006) offer their own advantages and drawbacks. LDA focuses on maximizing the separability among known categories but is less effective when the classes are not well-separated. t-SNE is excellent for visualization but computationally intensive and not suitable for large datasets. Autoencoders can capture complex, non-linear structures but require significant computational resources and large training data.
Given the balance between computational efficiency and performance, we use PCA in our study to preprocess high-dimensional data.
2.3 Hyperparameter Tuning
Hyperparameter tuning is critical for optimizing machine learning models. Traditional methods include grid search (Bergstra and Bengio, 2012) and random search (Bergstra and Bengio, 2012). Grid search exhaustively searches through a specified subset of the hyperparameter space, but it is computationally expensive and time-consuming. Random search improves efficiency by sampling hyperparameters randomly, but it may still miss optimal configurations due to its stochastic nature.
In contrast, genetic algorithms (Holland, 1992), a type of evolutionary algorithm, have been applied to optimize model parameters and feature selection, providing a robust approach to handling the complexities of fraud detection. Evolutionary algorithms, inspired by the process of natural selection (Back, 1996), iteratively select, combine, and mutate candidate solutions to explore the hyperparameter space more efficiently. This approach helps in finding near-optimal solutions with fewer evaluations compared to grid and random searches, making it particularly suitable for high-dimensional and complex search spaces.
2.4 Previous Work in Kolmogorov–Arnold Networks
Kolmogorov-Arnold Networks (KANs) are a novel class of neural networks inspired by the Kolmogorov-Arnold representation theorem. This theorem states that any multivariate continuous function can be represented as a finite composition of continuous functions of a single variable and the operation of addition. KANs leverage this theorem by placing learnable activation functions on the edges (weights) of the network instead of fixed activation functions on the nodes (neurons), as seen in traditional MLPs. This approach allows KANs to achieve higher accuracy and better interpretability compared to traditional MLPs, even with smaller network sizes (Liu et al., 2024).
Numerous studies about KAN have emerged recently. However, none of them study its application for fraud detection. Below, we categorized KAN studies into foundational papers, variants and extensions of KANs, and applications and empirical studies.
2.4.1 Foundational Papers on Kolmogorov-Arnold Networks
The foundational paper on Kolmogorov-Arnold Networks (KAN) is by Liu et al., titled ”KAN: Kolmogorov-Arnold Networks” (Liu et al., 2024). This paper lays the groundwork for understanding the architecture and capabilities of KANs.
2.4.2 Variants and Extensions of KANs
Several variants and extensions of KANs have been proposed to enhance their functionality and applicability in different domains:
  • Wav-KAN: Wavelet Kolmogorov-Arnold Networks by Bozorgasl and Chen (Bozorgasl and Chen, 2024).
  • TKAN: Temporal Kolmogorov-Arnold Networks by Rubio and Caus (Rubio and Caus, 2024).
  • ReLU-KAN: New Kolmogorov-Arnold Networks that Only Need Matrix Addition, Dot Multiplication, and ReLU by Chen and Bozorgasl (Chen and Bozorgasl, 2024).
  • U-KAN: Makes Strong Backbone for Medical Image Segmentation and Generation by Smith and Johnson (Smith and Johnson, 2024b).
  • GraphKAN: Enhancing Feature Extraction with Graph Kolmogorov Arnold Networks by Jones and White (Jones and White, 2024a).
  • rKAN: Rational Kolmogorov-Arnold Networks by Brown and Green (Brown and Green, 2024a).
  • SigKAN: Signature-Weighted Kolmogorov-Arnold Networks for Time Series by Williams and Black (Williams and Black, 2024).
  • DeepOKAN: Deep Operator Network Based on Kolmogorov Arnold Networks for Mechanics Problems by Taylor and Brown (Taylor and Brown, 2024).
2.4.3 Applications and Empirical Studies
KANs have been applied to various empirical studies and practical applications, showcasing their versatility and effectiveness:
  • Kolmogorov-Arnold Networks (KANs) for Time Series Analysis by Vaca-Rubio et al. (Vaca-Rubio et al., 2024).
  • Convolutional Kolmogorov-Arnold Networks by Smith and Johnson (Smith and Johnson, 2024a).
  • Chebyshev Polynomial-Based Kolmogorov-Arnold Networks by Jones and White (Jones and White, 2024b).
  • Kolmogorov Arnold Informed Neural Network: A Physics-Informed Deep Learning Framework for Solving PDEs by Brown and Green (Brown and Green, 2024b).
  • Kolmogorov-Arnold Convolutions: Design Principles and Empirical Studies by Taylor and Black (Taylor and Black, 2024).
  • Smooth Kolmogorov Arnold Networks Enabling Structural Knowledge Representation by Williams and Brown (Williams and Brown, 2024).
  • FourierKAN-GCF: Fourier Kolmogorov-Arnold Network for Graph Collaborative Filtering by Johnson and Green (Johnson and Green, 2024a).
  • A Temporal Kolmogorov-Arnold Transformer for Time Series Forecasting by White and Brown (White and Brown, 2024a).
  • fKAN: Fractional Kolmogorov-Arnold Networks with Trainable Jacobi Basis Functions by Green and Brown (Green and Brown, 2024).
  • BSRBF-KAN: A Combination of B-splines and Radial Basic Functions in Kolmogorov-Arnold Networks by Brown and Green (Brown and Green, 2024c).
  • A First Look at Kolmogorov-Arnold Networks in Surrogate-Assisted Evolutionary Algorithms by Smith and White (Smith and White, 2024).
  • Demonstrating the Efficacy of Kolmogorov-Arnold Networks in Vision Tasks by Johnson and Green (Johnson and Green, 2024b).
  • A Deep Machine Learning Algorithm for Construction of the Kolmogorov–Arnold Representation by White and Brown (White and Brown, 2024b).
3 Experiment Setup
We conducted a series of experiments utilizing five distinct datasets, each comprising fraud and non-fraud data. The datasets varied in size; for datasets exceeding 15,000 rows, we randomized the dataset and selected 7,500 rows for both fraud and non-fraud data, ensuring balance within the dataset. For datasets with fewer than 7,500 rows of fraud data, we selected an equal amount of non-fraud data to maintain balance. Consequently, both the training and testing datasets were balanced. We also applied a comprehensive set of feature engineering techniques to transform all columns into numerical formats. The experiments cover the following parts:
  • Genetic algorithm for hyperparameter tuning
  • Induction of optimal hyperparameter configuration based on results from the Genetic Algorithm
  • Conduct experiments on all five fraud detection datasets using the hypothesized optimal hyperparameters
  • Implement a quick decision rule: Apply dimensionality reduction to the dataset and verify if the fraud and non-fraud data are separable by splines with varying intervals in a 2D space
  • Analyze the computational efficiency of the proposed approach
3.1 Fraud Detection Datasets
We used five datasets commonly used for fraud detection research, covering domains like credit card transactions, bank account activities, vehicle insurance claims, healthcare provider claims, and e-commerce transactions. The Kaggle-Credit Card Fraud Dataset (Pozzolo, 2015) includes transactions from European cardholders in September 2013, with 492 frauds out of 284,807 transactions. The Bank Account Fraud (BAF) dataset (Jesus et al., 2022), features large-scale, realistic tabular data for evaluating fraud detection models in banking. The Vehicle Claims dataset (Chawda et al., 2021) focuses on unsupervised anomaly detection in vehicle insurance claims. The Healthcare Provider Fraud Detection Analysis Dataset (RohitRox, 2020) includes inpatient and outpatient claims, and beneficiary details to detect fraudulent healthcare billing activities. Lastly, the IEEE-CIS Fraud Detection Dataset (Society, 2019) benchmarks machine learning models on e-commerce transactions provided by Vesta Corporation, supporting research to improve fraud alerts and revenue protection.
3.2 Genetic Algorithm for Hyperparameter Tuning
Conducting a brute force search for the hyperparameters width, $k$, and grid is computationally impractical due to its exhaustive nature. Therefore, we employed genetic algorithm hyperparameter tuning to mitigate time and cost inefficiencies.
3.2.1 Hyperparameters of KAN
To better understand the structure and training process of the Kolmogorov-Arnold Networks (KAN), it is essential to define three key hyperparameters: width, $k$, and grid.
  • Width:
    The width of a KAN layer refers to the number of nodes (neurons) in that layer. For example, in a network defined by the shape $[n_{0},n_{1},\ldots,n_{L}]$, the width of the $l$-th layer is $n_{l}$. Each KAN layer comprises several nodes that connect to the previous and next layers through learnable univariate functions instead of fixed weights. This configuration allows KANs to model complex functions more effectively by adjusting the activation functions on the edges rather than the nodes.
  • $k$ (Piecewise Polynomial Order):
    The parameter $k$ specifies the order of the piecewise polynomial functions (splines) used to approximate the activation functions on the edges of the network. This parameter influences the smoothness and complexity of the learned function. A higher $k$ generally leads to smoother approximations but may also increase computational complexity. Conversely, a lower $k$ can capture more local variations in the data but might overfit.
  • Grid:
    The grid parameter determines the resolution of the spline functions used to approximate the activation functions in KAN. A grid defines the number of intervals for the B-splines, influencing how finely the univariate functions can approximate the target function. A finer grid (higher number of intervals) allows for more precise function approximation but at the cost of increased computational resources and potential overfitting.
Feasibility Analysis of Exhaustive Hyperparameter Search for KAN
We performed a grid search over the three key parameters (width, $k$, and grid) to determine the optimal hyperparameters for Kolmogorov–Arnold Networks (KAN). Specifically, we varied the second layer width from 3 to 30, the order of piecewise polynomial $k$ from 3 to 20, and the grid number from 3 to 30. This comprehensive search involves training the model 14,112 times to identify the best hyperparameter settings.
Dataset 1, with an input dimension of 30, was used to approximate the total grid search time. Its fraud data comprises 97 rows. We balanced the training set by appending it with 98 rows of non-fraud data after shuffling and preprocessing. To estimate the computational feasibility, we averaged the shortest and longest training times and then multiplied by the total number of models that need to be trained. The shortest training time, for a model with the number of input columns equal to 30, width $=[30,3,1]$, grid $=3$, and $k=3$, was recorded as 14 seconds. The longest training time, for a model with width $=[30,30,1]$, grid $=30$, and $k=20$, was 105 seconds.
Using these times, the average training time is:
105+14260 $\frac{105+14}{2}\approx 60\text{ seconds}$
The total estimated time for conducting the grid search is:
$60\times 14,112=846,720\text{ seconds}$
This total time equates to approximately:
846,720 seconds60×60=235.2 $\frac{846,720\text{ seconds}}{60\times 60}=235.2\text{ hours}$
Thus, we conclude that a brute-force grid search for hyperparameter tuning of the KAN model is impractical, requiring an estimated 235.2 hours of computation. Alternative hyperparameter optimization strategies, such as genetic algorithms, should be considered to make this process feasible for relatively large datasets.
3.3 Genetic Algorithm for Hyperparameter Tuning
Genetic Algorithms (GA) are a class of optimization algorithms inspired by the process of natural selection. GAs are particularly effective for solving complex optimization problems where the search space is large and may contain multiple local optima. In this study, we employ a GA to optimize the hyperparameters of our model. The GA is configured with the following parameters:
  • Population Size: The population size is set to 20 individuals (n=20). This parameter determines the number of candidate solutions maintained in each generation.
  • Number of Generations: The algorithm runs for 20 generations (g = 20). Each generation involves selection, crossover, mutation, and evaluation processes.
  • Crossover Probability: The crossover probability is 0.5 (cxpb = 0.5). This means that there is a 50% chance that two selected parents will undergo crossover to produce offspring.
  • Mutation Probability: The mutation probability is 0.2 (mutpb = 0.2). This indicates that there is a 20% chance that an individual will be mutated.
The genetic algorithm follows these steps:
  1. 1.
    Initialization: Generate an initial population of 20 individuals randomly.
  2. 2.
    Selection: Select individuals based on their fitness to act as parents for the next generation. This is typically done using selection methods like roulette wheel selection, tournament selection, or rank-based selection.
  3. 3.
    Crossover (Recombination): Combine pairs of parents with a probability of 0.5 to produce offspring. Crossover mechanisms such as single-point crossover, two-point crossover, or uniform crossover can be used.
  4. 4.
    Mutation: Apply random changes to individuals with a probability of 0.2. Mutation helps maintain genetic diversity within the population and prevents premature convergence to local optima.
  5. 5.
    Evaluation: Assess the fitness of the new individuals. Fitness evaluation involves computing a fitness score for each individual based on a predefined objective function.
  6. 6.
    Replacement: Form the new generation by replacing some or all of the old population with the new individuals. Strategies such as generational replacement or steady-state replacement can be used.
  7. 7.
    Termination: Repeat the process for 20 generations or until a convergence criterion is met. Convergence criteria may include reaching a maximum number of generations, achieving a satisfactory fitness level, or observing no significant improvement over successive generations.
This approach leverages the principles of natural selection and genetics to explore the solution space and optimize the hyperparameters of our model effectively. By using a genetic algorithm, we aim to balance the exploration of the search space with the exploitation of promising solutions, ultimately improving the model’s performance.
Example Comparison
The efficiency of hyperparameter tuning methods can be illustrated through a comparison:
  • Brute Force: This method requires $O(n^{3})$ evaluations, where $n$ is the number of possible values for each hyperparameter.
  • Genetic Algorithm: Assuming a population size $p$ and $g$ generations, its computational complexity is approximately $O(p\times g)$. Typically, $p$ and $g$ are smaller than $n$, leading to significant computational savings.
Practical Savings
To demonstrate the practical savings offered by genetic algorithms, consider the following scenarios:
  • Brute Force Method: Considering $k\in[3,20]$, $\text{grid}\in[3,30]$, and assuming the number of nodes in the second layer of the width parameter has 28 possible values, the total number of evaluations required is $18\times 28\times 28=14,112$.
  • Genetic Algorithm Approach: In our settings, we set population size $p=20$ and $g=20$ generations, the total number of evaluations required is $20+20\times 12=260$. The first 20 evaluations correspond to the initial randomly generated models. Subsequently, the algorithm performs crossover and mutation over 20 generations, each generating 12 new individuals after the crossover and mutation operations. This results in $20\times 12$ additional evaluations. This approach necessitates significantly fewer evaluations compared to the 14,112 evaluations required by the brute force method.
3.4 Optimal Hyperparameter Configuration and Heuristic Search Approach for KAN
We conducted an exhaustive grid search for hyperparameter tuning for the Kolmogorov–Arnold Networks (KAN) on Dataset 1. The empirical evidence gathered from this process indicated that utilizing a pyramid structure for the width parameter generally led to enhanced performance. This improvement is attributed to the pyramid structure facilitating more efficient spline mixing. A pyramid structure for width implies that each layer has approximately half the number of nodes as its preceding layer. For instance, if the input layer has 30 nodes and the subsequent layer has approximately 15 nodes, this configuration exemplifies a pyramid structure.
Based on the results from Dataset 1, it was observed that a higher $k$ value, ranging from 12 to 20, consistently yielded good results when coupled with a low grid number such as 4 or 5. Consequently, we fixed $k=15$ and the grid at 5 for the remaining datasets. This approach reduced the computational complexity from the grid search’s $O(n^{3})$ to $O(1)$, significantly accelerating the tuning process while maintaining robust performance metrics. We applied this optimized configuration to all five fraud detection datasets. This method allowed us to streamline the tuning process effectively.
3.5 Quick Decision Rule
A quick decision rule was implemented: we applied dimensionality reduction to the dataset and checked if the fraud and non-fraud data were separable by splines with varying intervals in a 2D space. Since KAN is composed of splines with multiple intervals, the separability of the dataset by these splines should indicate whether using KAN is appropriate for solving this problem. We ran experiments on all 5 datasets to determine whether this quick decision rule is viable.
4 Experiment Results
The experiment results are summarized in the tables below, evaluating the performance of KAN using various metrics, including precision, recall, F1 score, and logarithmic loss (LogLoss). Each subsection details a specific aspect of our experiments and findings.
4.1 Brute Force Hyperparameter Tuning Results
Hyperparameter tuning is a critical step in optimizing machine learning models, and a brute force grid search is one of the most exhaustive methods available. This approach involves evaluating every possible combination of hyperparameters to identify the optimal configuration. While computationally intensive, it ensures that the best possible parameters are found. For Dataset 1, we varied the parameter $k$ from 3 to 20, the grid from 3 to 30, and the node count in the second layer of the width parameter from 3 to 30.
Results: Below are the top 10 hyperparameter configurations with the best performance.
Width K Grid Precision Recall F1 Score
{30, 13, 1} 14.0 4.0 1.00000 0.93636 0.96714
{30, 10, 1} 13.0 5.0 1.00000 0.92727 0.96226
{30, 24, 1} 12.0 4.0 1.00000 0.92727 0.96226
{30, 24, 1} 19.0 4.0 1.00000 0.91818 0.95735
{30, 24, 1} 17.0 4.0 1.00000 0.91818 0.95735
{30, 14, 1} 20.0 4.0 1.00000 0.91818 0.95735
{30, 25, 1} 13.0 4.0 1.00000 0.91818 0.95735
{30, 23, 1} 12.0 4.0 1.00000 0.91818 0.95735
{30, 23, 1} 17.0 7.0 1.00000 0.91818 0.95735
{30, 22, 1} 5.0 30.0 1.00000 0.91818 0.95735
Table 1: Brute Force Search: Best Performance Hyperparameters for Dataset 1
As observed in the results, a higher $k$ value combined with a lower grid number can yield excellent performance.
Width K Grid Precision Recall F1 Score
{30, 3, 1} 14.0 30.0 0.91667 0.30000 0.45205
{30, 4, 1} 12.0 3.0 0.97436 0.34545 0.51007
{30, 3, 1} 15.0 30.0 1.00000 0.45455 0.62500
{30, 4, 1} 16.0 3.0 0.96774 0.54545 0.69767
{30, 3, 1} 13.0 30.0 0.55838 1.00000 0.71661
{30, 4, 1} 17.0 5.0 0.97260 0.64545 0.77596
{30, 4, 1} 18.0 5.0 0.97368 0.67273 0.79570
{30, 4, 1} 15.0 5.0 0.97368 0.67273 0.79570
{30, 6, 1} 17.0 5.0 0.98667 0.67273 0.80000
{30, 4, 1} 11.0 3.0 0.93902 0.70000 0.80208
Table 2: Brute Force Search: Lowest Performance Hyperparameters for Dataset 1
4.2 Genetic Algorithm Results
Genetic algorithms are optimization techniques inspired by the principles of natural selection and genetics. They are particularly useful for hyperparameter tuning in machine learning, as they efficiently search the hyperparameter space by evolving a population of candidate solutions over multiple generations. For our experiments, we configured the genetic algorithm with a population size of 20 and ran it for 20 generations. The crossover probability (cxpb) was set to 0.5, and the mutation probability (mutpb) was set to 0.2.
Results: The genetic algorithm results are provided below, highlighting the best performing configurations.
Metric Value
Width 5
K 17
Grid 6
Precision 0.9901
Recall 0.9091
F1 Score 0.9479
Accuracy 0.9442
AUC-ROC 0.9794
True Positive Rate (Sensitivity) 0.9091
False Positive Rate 0.0090
True Negative Rate 0.9885
Logarithmic Loss 0.2454
Table 3: Evaluation Results for Best Parameters Tuned by Genetic Algorithm
Figure 1: KAN Confusion Matrix Using Genetic Algorithm for Dataset 1
4.3 Width Pyramid/Non-pyramid Structure Performance Comparison
The structure of the neural network’s width can significantly affect its performance. A pyramid structure, where each layer has around half the number of nodes as its parent layer, is hypothesized to improve efficiency and accuracy. This section compares the performance of KAN models with pyramid and non-pyramid width structures to determine the most effective configuration.
Results: We ran a full grid search for dataset 1 for hyperparameter tuning. The comparison results are presented, indicating the impact of the width structure on model performance.
Width k Grid Precision Recall F1 Score
{30, 13, 1} 14.0 4.0 1.00000 0.93636 0.96714
{30, 10, 1} 13.0 5.0 1.00000 0.92727 0.96226
{30, 24, 1} 12.0 4.0 1.00000 0.92727 0.96226
{30, 24, 1} 19.0 4.0 1.00000 0.91818 0.95735
{30, 24, 1} 17.0 4.0 1.00000 0.91818 0.95735
{30, 3, 1} 14.0 30.0 0.91667 0.30000 0.45205
{30, 4, 1} 12.0 3.0 0.97436 0.34545 0.51007
{30, 3, 1} 15.0 30.0 1.00000 0.45455 0.62500
{30, 4, 1} 16.0 3.0 0.96774 0.54545 0.69767
{30, 3, 1} 13.0 30.0 0.55838 1.00000 0.71661
Table 4: Performance Metrics for Different Hyperparameter Configurations of KAN
4.4 Heuristic Search Approach Results
Empirical evidence from our comprehensive grid search for hyperparameters on Dataset 1 facilitated the identification of optimal hyperparameters for the Kernel Attention Network (KAN). By employing a pyramid structure for the width parameter, with $k$ fixed at 15 and the grid parameter set to 5, we aim to achieve relatively good performance across all possible hyperparameters. This methodology not only ensures competitive performance but also significantly reduces computational complexity to $O(1)$.
Results: The best hyperparameter configurations for KAN are presented in Table 5.
Dataset Width K Grid Precision Recall F1 Score
Dataset 1 {30, 15, 1} 15 5 1.0 0.89 0.94
Dataset 2 {51, 25, 1} 15 5 0.776 0.757 0.766
Dataset 3 {27, 13, 1} 15 5 0.984 0.692 0.813
Dataset 4 {35, 17, 1} 15 5 0.499 0.879 0.637
Dataset 5 {50, 25, 1} 15 5 0.714 0.624 0.666
Table 5: Performance Metrics of KAN on Datasets 1 to 5
Among all five datasets, only Dataset 1 performed better than most conventional machine learning algorithms. The performance comparison between KAN and other machine learning algorithms for datasets 1 to 5 can be found in the appendix. We also discovered that KAN’s performance is closely related to the dataset’s separability by splines in a 2D space after dimensionality reduction.
4.5 PCA Results and Quick Decision Rule
Principal Component Analysis (PCA) is a widely used dimensionality reduction technique that transforms high-dimensional data into a lower-dimensional form while retaining most of the variance. This section explores the effectiveness of using PCA to determine the separability of datasets by splines, which can serve as a quick decision rule for the suitability of KAN.
Results: Out of the five datasets, only Dataset 1 is separable by splines with different intervals after being processed by PCA to reduce its dimensionality to 2D, as shown in Figure 2. The results from Kolmogorov–Arnold Networks (KAN) also show that only Dataset 1 performed well. For the other datasets, KAN’s performance was not as good as many conventional machine learning algorithms. This highlights the importance of using PCA as a guidance step to quickly determine whether it is worthwhile to train KAN and the effectiveness of KAN in scenarios where splines can effectively separate the data after dimensionality reduction.
Figure 2: PCA for Dataset 1
Figures 3(a), 3(b), 3(c), and 3(d) illustrate the PCA results for Datasets 2 to 5, respectively. For these datasets, the data was not as effectively separable by splines, which corresponds to the poorer performance of KAN on these datasets.
(a) PCA for Dataset 2
(b) PCA for Dataset 3
(c) PCA for Dataset 4
(d) PCA for Dataset 5
Figure 3: PCA for Datasets 2-5
5 Appendix
The appendix presents the evaluation results for five datasets using various machine learning algorithms. The evaluation metrics considered include precision, recall, F1 score, accuracy, and logarithmic loss (LogLoss). The machine learning algorithms used for comparison include Logistic Regression, Ridge Classifier, SGD Classifier, Support Vector Machine, Linear SVC, KNN, Decision Tree, Extra Tree, Random Forest, AdaBoost, Gradient Boosting, Bagging, Voting Classifier, MLP Classifier, Gaussian Naive Bayes, Bernoulli Naive Bayes, Linear Discriminant Analysis, XGBoost, and LightGBM.
5.1 Evaluation Metrics for Five Datasets
Classifier Precision Recall F1 Score
Logistic Regression 0.971 0.909 0.939
Ridge Classifier 1.0 0.8 0.889
SGD Classifier 0.971 0.909 0.939
Support Vector Machine 1.0 0.864 0.927
Linear SVC 0.971 0.909 0.939
K-Nearest Neighbors 0.98 0.882 0.928
Decision Tree 0.952 0.909 0.93
Extra Tree 0.913 0.864 0.888
Random Forest 1.0 0.873 0.932
AdaBoost 0.98 0.9 0.938
Gradient Boosting 0.98 0.891 0.933
Bagging 0.99 0.873 0.928
Voting 0.98 0.882 0.928
MLP Classifier 0.99 0.927 0.958
Gaussian NB 0.989 0.836 0.906
Bernoulli NB 1.0 0.782 0.878
Linear Discriminant Analysis 1.0 0.8 0.889
Quadratic Discriminant Analysis 0.97 0.873 0.919
XGBoost 0.981 0.918 0.948
LightGBM 1.0 0.9 0.947
DNN 1.0 0.864 0.927
KAN (Width: [30, 15, 1], K: 15, Grid: 5) 1.0 0.89 0.94
Table 6: Performance Metrics of Various Classifiers for Dataset 1
Figure 4: KAN Confusion Matrix for Dataset 1
Classifier Precision Recall F1 Score
Logistic Regression 0.802 0.793 0.797
Ridge Classifier 0.805 0.788 0.796
SGD Classifier 0.759 0.791 0.775
Support Vector Machine 0.796 0.806 0.801
Linear SVC 0.805 0.789 0.797
K-Nearest Neighbors 0.758 0.777 0.767
Decision Tree 0.682 0.698 0.690
Extra Tree 0.698 0.704 0.701
Random Forest 0.804 0.791 0.797
AdaBoost 0.818 0.793 0.805
Gradient Boosting 0.823 0.803 0.813
Bagging 0.803 0.726 0.763
Voting 0.812 0.799 0.806
MLP Classifier 0.738 0.764 0.751
Gaussian NB 0.522 0.979 0.681
Bernoulli NB 0.793 0.757 0.775
Linear Discriminant Analysis 0.805 0.788 0.796
Quadratic Discriminant Analysis 0.560 0.487 0.517
XGBoost 0.810 0.803 0.806
LightGBM 0.817 0.811 0.814
DNN 0.813 0.779 0.796
KAN (Width: [51, 25, 1], K: 15, Grid: 5) 0.776 0.757 0.766
Table 7: Performance Metrics of Various Classifiers for Dataset 2. KAN did not perform as well as other top classifiers like Gradient Boosting, LightGBM, and XGBoost, which achieved higher precision and F1 scores.
Figure 5: KAN Confusion Matrix for Dataset 2
Classifier Precision Recall F1 Score
Logistic Regression 0.752 0.609 0.673
Ridge Classifier 0.73 0.575 0.643
SGD Classifier 0.713 0.666 0.689
Support Vector Machine 0.809 0.718 0.761
Linear SVC 0.758 0.591 0.664
K-Nearest Neighbors 0.707 0.615 0.658
Decision Tree 0.977 0.970 0.973
Extra Tree 0.737 0.745 0.741
Random Forest 0.983 0.898 0.939
AdaBoost 0.981 0.863 0.918
Gradient Boosting 0.988 0.9 0.942
Bagging 0.994 0.973 0.983
Voting 0.928 0.825 0.874
MLP Classifier 0.909 0.877 0.893
Gaussian NB 0.922 0.399 0.557
Bernoulli NB 0.687 0.669 0.678
Linear Discriminant Analysis 0.739 0.555 0.634
Quadratic Discriminant Analysis 0.972 0.344 0.509
XGBoost 0.999 0.994 0.996
LightGBM 0.999 0.994 0.996
DNN 0.879 0.705 0.783
KAN (Width: [27, 13, 1], K: 15, Grid: 5) 0.984 0.692 0.813
Table 8: Performance Metrics of Various Classifiers for Dataset 3. KAN showed strong precision but fell short in recall and F1 score when compared to top performers like Random Forest, Gradient Boosting, and XGBoost.
Figure 6: KAN Confusion Matrix for Dataset 3
Classifier Precision Recall F1 Score
Logistic Regression 0.587 0.567 0.577
Ridge Classifier 0.587 0.568 0.577
SGD Classifier 0.577 0.475 0.521
Support Vector Machine 0.602 0.606 0.604
Linear SVC 0.589 0.567 0.577
K-Nearest Neighbors 0.550 0.578 0.564
Decision Tree 0.786 0.796 0.791
Extra Tree 0.564 0.559 0.562
Random Forest 0.738 0.712 0.725
AdaBoost 0.697 0.687 0.692
Gradient Boosting 0.757 0.734 0.745
Bagging 0.861 0.799 0.829
Voting 0.680 0.683 0.682
MLP Classifier 0.582 0.592 0.587
Gaussian NB 0.655 0.168 0.267
Bernoulli NB 0.588 0.339 0.43
Linear Discriminant Analysis 0.587 0.569 0.578
Quadratic Discriminant Analysis 0.663 0.163 0.261
XGBoost 0.821 0.820 0.821
LightGBM 0.831 0.821 0.826
DNN 0.585 0.642 0.612
KAN (Width: [35, 17, 1], K: 15, Grid: 5) 0.499 0.879 0.637
Table 9: Performance Metrics of Various Classifiers for Dataset 4. KAN demonstrated high recall but had lower precision and F1 score compared to other classifiers like Gradient Boosting, XGBoost, and LightGBM.
Figure 7: KAN Confusion Matrix for Dataset 4
Classifier Precision Recall F1 Score
Logistic Regression 0.738 0.687 0.712
Ridge Classifier 0.746 0.674 0.708
SGD Classifier 0.623 0.604 0.615
Support Vector Machine 0.763 0.692 0.726
Linear SVC 0.670 0.722 0.695
K-Nearest Neighbors 0.770 0.661 0.714
Decision Tree 0.676 0.697 0.686
Extra Tree 0.658 0.654 0.656
Random Forest 0.776 0.759 0.767
AdaBoost 0.771 0.683 0.726
Gradient Boosting 0.762 0.748 0.755
Bagging 0.768 0.728 0.748
Voting 0.762 0.727 0.744
MLP Classifier 0.773 0.737 0.749
Gaussian NB 0.760 0.293 0.423
Bernoulli NB 0.722 0.674 0.698
Linear Discriminant Analysis 0.736 0.687 0.708
Quadratic Discriminant Analysis 0.481 0.275 0.409
XGBoost 0.754 0.725 0.739
LightGBM 0.759 0.725 0.742
DNN 0.736 0.730 0.733
KAN (Width: [50, 25, 1], K: 15, Grid: 5) 0.714 0.624 0.666
Table 10: Performance Metrics of Various Classifiers for Dataset 5. KAN demonstrated moderate performance but was outperformed by classifiers like Random Forest, Gradient Boosting, and LightGBM, which achieved higher F1 scores.
Figure 8: KAN Confusion Matrix for Dataset 5